please with the steps!


NH4NO2(s) → N2(g) + H2O(g)



When a sample is decomposed in a test tube, 747 mL of wet N2(g) is collected over water at 27°C and 558 torr total pressure. How many grams of dry NH4NO2(s) were initially decomposed? The vapor pressure of water at 27°C is 24.3 torr.


Mass ? in grams




1
Expert's answer
2021-03-10T06:16:43-0500

1.find the amount of substance

"PV = nRT"

P=558-24,3=533,7=0.702237 atm

27°C=300,15 K

"0.702237\times0.747= n\times0.08206\times300.15"

"0.52 = n\times24.63"

n=0.0211

"m=M\times n=64.044\times0.0211=1.35 g"